# An agent-based model of behaviour change calibrated to reversal learning data

### v1

@@ Abstract
Behaviour change lies at the heart of many observable collective phenomena such as the transmission and control of infectious diseases, adoption of public health policies, and migration of animals to new habitats. Representing the process of individual behaviour change in computer simulations of these phenomena remains an open challenge. Often, computational models use phenomenological implementations with limited support from behavioural data. Without a strong connection to observable quantities, such models have limited utility for simulating observed and counterfactual scenarios of emergent phenomena because they cannot be validated or calibrated. Here, we present a simple stochastic individual-based model of reversal learning that captures fundamental properties of individual behaviour change, namely, the capacity to learn based on accumulated reward signals, and the transient persistence of learned behaviour after rewards are removed or altered. The model has only two parameters, and we use approximate Bayesian computation to demonstrate that they are fully identifiable from empirical reversal learning time series data. Finally, we demonstrate how the model can be extended to account for the increased complexity of behavioural dynamics over longer time scales involving fluctuating stimuli. This work is a step towards the development and evaluation of fully identifiable individual-level behaviour change models that can function as validated submodels for complex simulations of collective behaviour change.
individual-based model, agent-based model, behavioural change, reversal learning, calibration, approximate Bayesian computation
@@ I Introduction
Behavioural changes are fundamental to many real-world phenomena we seek to understand or predict through computer simulations. Such computational models describe how the actions of individuals in a system change in response to events caused by internal or external factors. Variation in behaviour over time affects the dynamics of the modelled real-world systems and the phenomena that emerge from them, from how infectious diseases spread in a population to how animals adapt to new habitats Verelst2016 ; Weston2018 ; Bauer2013 ; Leimar2024 . However, in many computer simulations, these individual-level submodels are based on strong assumptions about the underlying factors that influence behaviour and the conditions that drive behaviour change, with limited empirical support. Hence, such models cannot be validated or calibrated based on behavioural data, raising concerns about the verifiability of their results Verelst2016 ; Weston2018 . There remains a need for explicit computational representations of behaviour change that are supported by quantitative empirical evidence while maintaining the mechanistic adequacy to simulate observed and counterfactual scenarios. To help address this need, we present a simple (two-parameter) stochastic model of individual behaviour change and demonstrate that its parameters are identifiable from empirical time series data from reversal learning experiments.
@@
To introduce our model, we begin with the general premise that individual decisions about behaviour map available information (stimuli) to a set of actions (responses). As a motivating example, consider endogenous protective behaviours during an infectious disease outbreak: The choice of some individuals to adopt social distancing practices depends on external information from their environment which could be globally available from (e.g.) government and media reports of disease statistics Collinson2015 ; Xiao2015 . Additional local information may also be available from communication and observations involving neighbours, friends, and family Andrews2015 . Such external information can support or contradict information internal to individuals, such as their perceived risks and benefits. How individuals then incorporate this information into their decision-making process may vary from person to person. Such variation often correlates with observable factors such as age, occupation, or socioeconomic status, which may reflect different capabilities and considerations that individuals have when making decisions Gozzi2021 ; Gauvin2021 ; Zachreson2021 . Thus, when modelling behaviour in computer simulations, design choices are often made about the types of information individuals can use to make decisions and the mechanisms by which they translate that information into actions (see, e.g., Durham2012 ).
@@
In behaviour change models, environmental stimulus could be the number of infected people in an infectious disease model Poletti2011 ; Durham2012 ; Bayham2015 , the physical features of prey in an animal foraging model Dyer2014 ; Leimar2024 , or a reward indicating the desirability of a chosen action from the perspective of a global observer with complete knowledge Xia2021 ; Le2023 . In simulations, there are two general approaches to modelling how individuals act on this information: rule-based or learning-based. Rule-based approaches condition an individual’s behaviour on external information based on predefined threshold criteria or logical relationships Durham2012 ; Bayham2015 . In contrast, learning-based approaches assume that individuals are rational and behave optimally by choosing actions that maximize an objective function. The objective functions can be elaborate and may incorporate computing the costs and benefits of multiple strategies or examining past experiences Poletti2011 ; Dyer2014 .
@@
All studies in which such models are developed must make assumptions on the factors that drive behaviour change, but many do so without being calibrated or validated from empirical data Verelst2016 ; Weston2018 . In those studies that use behavioural data, the information is often in the form of surveys which are costly to conduct and typically quantify behavioural intentions rather than actual behaviours. Additionally, most previous modelling studies do not draw from the rich body of literature that examines and describes the psychology behind behaviour change which has the potential to provide a stronger theoretical foundation for computational models Weston2020 . By designing a computational representation of behaviour that is informed by both experimental data and behavioural change theories, we can improve the validity, reliability, and practicality of computer simulations.
@@
To develop our individual-based model of behaviour change, we aim to simulate the reversal learning experiment that is often conducted in neuroscience, psychology, and ecology to study how individuals adapt their behaviour to changes in their environment. An overview of the reversal learning experiment setup is shown in Figure 1. In a typical reversal learning experiment, an individual is tasked to choose between two options: a rewarded and a non-rewarded option. The experiment is repeated in a series of trials and examines whether the individual can learn which of the two options is rewarded. After some time, the reward association is reversed. The individual’s response to the change in the stimulus is then recorded over a second set of repeated trials over which the new reward association may be learned Rayburn-reeves2011 ; Dyer2014 ; Hassett2017 ; Xia2021 ; Le2023 ; Leimar2024 .
By simulating reversal learning, we developed an agent-based model of behaviour change with parameters that are fully identifiable from empirical data. We show that this model satisfies three characteristics of a good behavioural change model: (1) mechanistic adequacy, (2) tangibility, and (3) tractability. By mechanistic adequacy, we refer to the degree to which the model’s underlying processes are relatable to the features of the behaviour change process (in this case those characterising reversal learning). Tangibility means the model is tangible: its parameters have intuitive meaning and it produces output that can be directly compared to observable quantities or patterns. Tractability means the model is tractable: its parameters are identifiable based on comparison with experiments and it can be calibrated efficiently in practice with sufficiently low computational cost.
@@ II Methods
In this section, we discuss how our agent-based model represents reversal learning behaviour. We then describe the experimental data we used to understand and evaluate this behaviour and the experiments that generated the data. Finally, we describe how calibration was performed.
@@ II.1 Model description
We developed a simple agent-based model of behaviour change where an agent’s behaviour mimics that which is observed in reversal learning experiments. An overview of our agent-based model is shown in Figure 2. Abstractly, agents represent any individual capable of making decisions that could produce reward signals. In the model, the agent decides on an action $a$ from a given set of possible actions at every iteration of the simulation. We designed the agent to be capable of sensing its environment and remembering its past experiences which enable it to acquire and learn information for deciding which action to choose, resembling individuals in the real world. The information from the environment comes in the form of a stimulus $s$ that encapsulates the impact of the agent’s action on the environment in terms of an associated reward signal. The information internal to the agent is encoded as a memory value $m$ that quantifies the desirability of an action based on the agent’s past experiences. Given this information, the agent decides whether to stay with or change its current action based on the probability returned by its logistic decision function. The agent’s behavioural tendencies are characterized by the timescale of its memory $\tau$, and the logistic decision function’s steepness $k$ which specifies the agent’s certainty about the value of the information it has accumulated. In every simulation, the environment periodically switches the stimulus associated with the agent’s actions and observes the agent’s probability of choosing the rewarded or non-rewarded actions over time. Below, we provide a full model description following the ODD (Overview, Design concepts, Details) protocol for describing individual- and agent-based models Grimm2006 , as updated by Grimm2020 .
@@
Figure 2: Schematic of the decision-making process in the agent-based model. (a) The agent chooses an action $a_{t}$ at time step $t$ and receives a corresponding stimulus $s_{t}$ from the environment. (b) In choosing the next action $a_{t+1}$, the agent considers the utility $u_{t}$ of its current action $a_{t}$. This utility is the sum of the instantaneous stimulus $s_{t}$ from the environment and the agent’s internal memory $m_{t}$ of past actions and stimuli with a timescale determined by the memory decay $\tau$. The agent decides the next action $a_{t+1}$ using a logistic function with a steepness $k$, which indicates the agent’s certainty on the informational value of any given utility.
II.1.1 Purpose and patterns
The model’s specific purpose is to illustrate a simple computational representation of a single agent’s behaviour that can be calibrated and validated from empirical data. Having such a representation of behaviour that is mechanistically adequate while being empirically identifiable will enable the incorporation of validated behavioural submodels into simulations of multi-agent systems. The patterns used as the criterion for evaluating the model’s suitability for its purpose are the time series of an agent’s decisions (for an individual) and the time series of the success rate (for an ensemble).
II.1.2 Entities, state variables, and scales
There are two entities in the model. The first entity is the environment which facilitates the reversal learning experiment. The state variables of the environment are the list of possible actions that the agent can choose from, the stimulus $s$ that is given to the agent for each action, the number of trial periods $n$, and the number of trials $r$ in each trial period. All the environment’s state variables can be specified from the setup of empirical experiments and are static variables except for the stimulus $s$ associated with each action which changes per trial period.
The second entity in the model is the agent whose behaviour is observed during the experiment. The state variables of the agent are its chosen action $a$ at every time step, the corresponding stimulus $s$ that it receives from the environment, a memory value $m$ which represents the agent’s preference for the chosen action $a$ based on its memory of past experiences (with timescale determined by the memory decay $\tau$), and a logistic function with steepness $k$ which represents the agent’s decision function. The action $a$, stimulus $s$, and memory $m$ are dynamic variables that change over time, while the memory decay $\tau$ and steepness $k$ are static parameters that characterize an agent’s behavioural tendencies.
Space is not represented in the model. Time is represented in discrete increments with an abstract unit of 1 trial and an extent equal to the number of trials $r$ multiplied by the number of trial periods $n$ in a given reversal learning experiment. This representation of time follows the setup used in the empirical experiments to which we refer for calibration and validation Buechel2018 ; Boussard2020 .
II.1.3 Process overview and scheduling
At every time step $t$, the model runs a simple 2-step process:
  1. 1.
    The agent chooses an action $a_{t}$ by executing the Choose_action submodel.
  2. 2.
    The environment gives the agent a stimulus $s_{t}$ associated with the agent’s chosen action $a_{t}$. The stimulus $s_{t}$ associated with each action changes per trial period.
II.1.4 Design concepts
The design concepts implemented in the agent-based model are basic principles, emergence, adaptation, learning, sensing, stochasticity, and observation. The agent-based model includes no objectives, prediction, and collectives.
Basic principles. The model is designed to demonstrate how behaviour could be represented in computer simulations such that it can be calibrated and validated from observable behavioural data. Moreover, the components we included in our computational model are informed broadly by the COM-B model of behaviour change which is a widely used theoretical framework on what drives behavioural changes (see Discussion for more information).
Emergence. The main result from the model which is the observed success rate over time emerges from the stimuli that the agent receives from the environment as well as its memory of past actions and stimuli. This probability of choosing an action is also partly imposed by the agent’s parameters for the memory decay $\tau$ and the steepness $k$ of the decision function which specifies the behavioural tendencies of the agent.
Adaptation. The agent has one adaptive behavior which is to choose an action at every time step. The agent’s decision of which action to choose is a probabilistic process where the probabilities are determined by a logistic function that takes as input the sum of the instantaneous stimulus that the agent receives from the environment and the agent’s internal memory of past experiences.
Learning. The adaptive behaviour of the agent incorporates learning through the use of an internal memory that stores the agent’s past actions and their associated stimuli. This memory enables the agent to learn the association between actions and stimuli and is developed as the agent gets more experience. The influence of past actions and stimuli on the agent’s decision at every time step decays exponentially over time. This learning process is detailed in the Get_memory submodel (described in Section 2).
Sensing. The agent is assumed to sense exactly the instantaneous stimulus associated with its action (it knows whether or not it made the rewarded decision). This sensing mechanism provides the agent with immediate feedback on the desirability of its most recent action.
Stochasticity. The action that the agent chooses at each time step is designed as a stochastic process to represent non-deterministic behaviour.
Observation. The fraction of agents choosing the rewarded action over time is observed from the relative frequency with which the agent chooses the action across an ensemble of simulation runs.
II.1.5 Initialization
At initialization, the list of possible actions that the agent can choose from and their associated stimulus are specified. Here, the agent can choose between options A and B. One of these options is initially assigned with a positive stimulus of +1+1, while the other has a stimulus of 0 (no stimulus). These stimulus values were decided based on the experimental setup in the studies we used for validation and calibration (see Section II.2).
II.1.6 Input data
The model does not use input data to represent time-varying processes. However, the model could be extended to allow a time-varying stimulus based on time series input.
II.1.7 Submodels
The model has 2 submodels: (1) the Get_memory submodel which details the agent’s memory process and (2) the Choose_action submodel which describes the mechanics of the agent’s decision-making process.
@@ II.1.7 Submodels
Get_memory submodel. We assume that the agent’s past actions and stimuli have an exponentially decaying influence on the agent’s memory. If we consider some previous time step $i=t-\Delta t$, the influence $w_{i}$ of the experience at the previous time step $i$ to the memory $m_{t}$ at the current time step $t$ is expressed as:
@@
where $a_{t}$ is the agent’s action at the current time step $t$, $a_{i}$ is the agent’s action at the previous time step $i$, $s_{i}$ is the corresponding stimulus that the agent received for action $a_{i}$, $\Delta t$ is the elapsed time between time steps $t$ and $i$, and $\tau$ is the timescale of memory. A higher $\tau$ indicates longer memory, while a smaller $\tau$ means shorter memory. The memory $m_{t}$ of all past actions and stimuli is then computed as:
$m_{t}=\sum_{i=1}^{t-1}{w_{i}}$ (2)
@@
We excluded the current experience at time step $t$ from the memory calculation to enable the current stimulus $s_{t}$ to be weighed differently from the memory $m_{t}$ of past actions and stimuli. Moreover, note that the sign of the influence $w$ in Equation 1 is relative to the current action $a_{t}$. Hence, positive values influence the agent to stay with the current action $a_{t}$ while negative values influence the agent to change the current action $a_{t}$ in the next time step $t+1$.
Choose_action submodel. The agent chooses its next action $a_{t+1}$ at the next time step $t+1$ based on the utility of its current action $a_{t}$ at the current time step $t$. The agent receives a stimulus $s_{t}$ from the environment corresponding to its current action $a_{t}$. The agent also computes a memory value $m_{t}$ for the current action $a_{t}$ using the Get_memory submodel. This memory value indicates the agent’s preference for the current action $a_{t}$ based on its past experiences. The utility $u_{t}$ of the current action $a_{t}$ is then defined as the sum of the environmental stimulus $s_{t}$ and the agent’s internal memory value $m_{t}$:
$u_{t}=s_{t}+m_{t}$ (3)
This utility represents the combination of all factors that the agent considers when making a decision. The utility $u_{t}$ is then passed to a logistic decision function:
$p_{a_{t+1}}=\frac{1}{1+e^{-ku_{t}}}$ (4)
@@
where the steepness $k$ indicates the agent’s certainty on the informational value of any given utility. The logistic decision function returns a probability $p_{a_{t+1}}$ indicating the agent’s likelihood of staying with the current action $a_{t}$ in the next time step $t+1$. Hence, the agent’s next action $a_{t+1}$ will be the same as the current action $a_{t}$ with probability $p_{a_{t+1}}$ or will be the other action with probability $1-p_{a_{t+1}}$.
II.2 Reversal learning experiments

### v2

# A microsimulation model of behaviour change calibrated to reversal learning data

@@ Abstract
Behaviour change lies at the heart of many observable collective phenomena such as the transmission and control of infectious diseases, adoption of public health policies, and migration of animals to new habitats. Representing the process of individual behaviour change in computer simulations of these phenomena remains an open challenge. Often, computational models use phenomenological implementations with limited support from behavioural data. Without a strong connection to observable quantities, such models have limited utility for simulating observed and counterfactual scenarios of emergent phenomena because they cannot be validated or calibrated. Here, we present a simple stochastic microsimulation model of reversal learning that captures fundamental properties of individual behaviour change, namely, the capacity to learn based on accumulated reward signals, and the transient persistence of learned behaviour after rewards are removed or altered. The model has only two parameters, and we use approximate Bayesian computation to demonstrate that they are fully identifiable from empirical reversal learning time series data. Finally, we demonstrate how the model can be extended to account for the increased complexity of behavioural dynamics over longer time scales involving fluctuating stimuli. This work is a step towards the development and evaluation of fully identifiable individual-level behaviour change models that can function as validated submodels for complex simulations of collective behaviour change.
microsimulation, behavioural change, reversal learning, calibration, approximate Bayesian computation
@@ I Introduction
Behavioural changes are fundamental to many real-world phenomena we seek to understand or predict through computer simulations. Such computational models describe how the actions of individuals in a system change in response to events caused by internal or external factors. Variation in behaviour over time affects the dynamics of the modelled real-world systems and the phenomena that emerge from them, from how infectious diseases spread in a population to how animals adapt to new habitats Verelst et al., (2016); Weston et al., (2018); Bauer and Klaassen, (2013); Leimar et al., (2024). However, in many computer simulations, these individual-level submodels are based on strong assumptions about the underlying factors that influence behaviour and the conditions that drive behaviour change, with limited empirical support. Hence, such models cannot be validated or calibrated based on behavioural data, raising concerns about the verifiability of their results Verelst et al., (2016); Weston et al., (2018). There remains a need for explicit computational representations of behaviour change that are supported by quantitative empirical evidence while maintaining the mechanistic adequacy to simulate observed and counterfactual scenarios. To help address this need, we present a simple (two-parameter) stochastic model of individual behaviour change and demonstrate that its parameters are identifiable from empirical time series data from reversal learning experiments.
@@
To introduce our model, we begin with the general premise that individual decisions about behaviour map available information (stimuli) to a set of actions (responses). As a motivating example, consider endogenous protective behaviours during an infectious disease outbreak: The choice of some individuals to adopt social distancing practices depends on external information from their environment which could be globally available from (e.g.) government and media reports of disease statistics Collinson et al., (2015); Xiao et al., (2015). Additional local information may also be available from communication and observations involving neighbours, friends, and family Andrews and Bauch, (2015). Such external information can support or contradict information internal to individuals, such as their perceived risks and benefits. How individuals then incorporate this information into their decision-making process may vary from person to person. Such variation often correlates with observable factors such as age, occupation, or socioeconomic status, which may reflect different capabilities and considerations that individuals have when making decisions Gozzi et al., (2021); Gauvin et al., (2021); Zachreson et al., (2021). Thus, when modelling behaviour in computer simulations, design choices are often made about the types of information individuals can use to make decisions and the mechanisms by which they translate that information into actions (see, e.g., Durham and Casman, (2012)).
@@
In behaviour change models, environmental stimulus could be the number of infected people in an infectious disease model Poletti et al., (2011); Durham and Casman, (2012); Bayham et al., (2015), the physical features of prey in an animal foraging model Dyer et al., (2014); Leimar et al., (2024), or a reward indicating the desirability of a chosen action from the perspective of a global observer with complete knowledge Xia et al., (2021); Le et al., (2023). In simulations, there are two general approaches to modelling how individuals act on this information: rule-based or learning-based. Rule-based approaches condition an individual’s behaviour on external information based on predefined threshold criteria or logical relationships Durham and Casman, (2012); Bayham et al., (2015). In contrast, learning-based approaches assume that individuals are rational and behave optimally by choosing actions that maximize an objective function. The objective functions can be elaborate and may incorporate computing the costs and benefits of multiple strategies or examining past experiences Poletti et al., (2011); Dyer et al., (2014).
@@
All studies in which such models are developed must make assumptions on the factors that drive behaviour change, but many do so without being calibrated or validated from empirical data Verelst et al., (2016); Weston et al., (2018). In those studies that use behavioural data, the information is often in the form of surveys which are costly to conduct and typically quantify behavioural intentions rather than actual behaviours. Additionally, most previous modelling studies do not draw from the rich body of literature that examines and describes the psychology behind behaviour change which has the potential to provide a stronger theoretical foundation for computational models Weston et al., (2020). By designing a computational representation of behaviour that is informed by both experimental data and behavioural change theories, we can improve the validity, reliability, and practicality of computer simulations.
To develop our microsimulation model of behaviour change, we aim to simulate the reversal learning experiment that is often conducted in neuroscience, psychology, and ecology to study how individuals adapt their behaviour to changes in their environment. An overview of the reversal learning experiment setup is shown in Figure 1. In a typical reversal learning experiment, an individual is tasked to choose between two options: a rewarded and a non-rewarded option. The experiment is repeated in a series of trials and examines whether the individual can learn which of the two options is rewarded. After some time, the reward association is reversed, but the individual is unaware of this reversal. The individual’s response to this change in stimulus is then recorded over a second set of repeated trials over which the new reward association may be learned Rayburn-Reeves et al., (2011); Dyer et al., (2014); Hassett and Hampton, (2017); Xia et al., (2021); Le et al., (2023); Leimar et al., (2024).
@@
By simulating reversal learning, we developed a microsimulation model of behaviour change with parameters that are fully identifiable from actual behavioural data. To identify the values of these parameters, we used approximate Bayesian computation which is an established approach for calibrating individual-level models to empirical data Van Der Vaart et al., (2015); Grazzini et al., (2017). We show that this model satisfies three characteristics of a good behavioural change model: (1) mechanistic adequacy, (2) tangibility, and (3) tractability. By mechanistic adequacy, we refer to the degree to which the model’s underlying processes are relatable to the features of the behaviour change process (in this case those characterising reversal learning). Tangibility means the model is tangible: its parameters have intuitive meaning and it produces output that can be directly compared to observable quantities or patterns. Tractability means the model is tractable: its parameters are identifiable based on comparison with experiments and it can be calibrated efficiently in practice with sufficiently low computational cost. In addition, we explain how our microsimulation model relates to the COM-B model of behaviour change, a theoretical framework describing that capability (C), opportunity (O), and motivation (M) are the three components necessary for a behaviour (B) to occur.
@@ II Methods
In this section, we discuss how our microsimulation model represents reversal learning behaviour. We then describe the experimental data we used to understand and evaluate this behaviour and the experiments that generated the data. Finally, we describe how calibration was performed.
@@ II.1 Model description
We developed a simple microsimulation model of behaviour change where an individual’s behaviour mimics that which is observed in reversal learning experiments. An overview of our microsimulation model is shown in Figure 2. Abstractly, individuals represent any entity capable of making decisions that could produce reward signals. In the model, the individual decides on an action $a$ from a given set of possible actions at every iteration of the simulation. Here, we consider only binary choices for simplicity, but the model can be extended to relax this constraint (see Discussion).
@@
Figure 2: Schematic of the decision-making process in the microsimulation model. (a) The individual chooses an action $a_{t}$ at time step $t$ and receives a corresponding stimulus $s_{t}$ from the environment. (b) In choosing the next action $a_{t+1}$, the individual considers the utility $u_{t}$ of its current action $a_{t}$. This utility is the sum of the instantaneous stimulus $s_{t}$ from the environment and the individual’s internal memory $m_{t}$ of past actions and stimuli with a timescale determined by the memory decay $\tau$. The individual decides the next action $a_{t+1}$ using a logistic function with a steepness $k$, which indicates the individual’s certainty on the informational value of any given utility.
We created the individual to be capable of sensing its environment and remembering its past experiences which enable it to acquire and learn information for deciding which action to choose, resembling decision-making entities in the real world. Suppose that at the current time step $t$, the individual chooses action $a_{t}$. The information from the environment comes in the form of a stimulus $s_{t}$ that encapsulates the impact of the individual’s action $a_{t}$ on the environment. How the environment provides this stimulus $s_{t}$ depends on the specific context the model is used for. In the context of reversal learning experiments, we set it up as:
[…]
In every simulation of the reversal learning experiment, the environment periodically switches the stimulus associated with the individual’s actions and observes the individual’s probability of choosing the rewarded actions over time.
@@
The information internal to the individual is encoded as a memory value $m_{t}$ that quantifies the desirability of an action $a_{t}$ based on the individual’s past experiences. We assume that the individual’s past actions and stimuli have an exponentially decaying influence on the individual’s memory. If we consider some previous time step $i=t-\Delta t$, the influence $w_{i}$ of the experience at the previous time step $i$ to the memory $m_{t}$ at the current time step $t$ is expressed as:
@@
where $a_{t}$ is the individual’s action at the current time step $t$, $a_{i}$ is the individual’s action at the previous time step $i$, $s_{i}$ is the corresponding stimulus that the individual received for action $a_{i}$, $\Delta t$ is the elapsed time between time steps $t$ and $i$, and $\tau$ is the timescale of memory. Note that the sign of the influence $w_{i}$ in Equation 2 is relative to the current action $a_{t}$. Hence, positive values influence the individual to stay with the current action $a_{t}$ while negative values influence the individual to change the current action $a_{t}$ in the next time step $t+1$. The memory $m_{t}$ of all past actions and stimuli is then computed as:
$m_{t}=\sum_{i=1}^{t-1}{w_{i}}.$ (3)
We note that the current experience at time step $t$ is excluded from the memory calculation to enable the instantaneous stimulus $s_{t}$ to be weighed differently from the memory $m_{t}$ of past actions and stimuli.
Together, the external stimulus $s_{t}$ from the environment and the internal memory $m_{t}$ of the individual describe the utility $u_{t}$ of the individual’s current action $a_{t}$ at the current time step $t$:
$u_{t}=s_{t}+m_{t}.$ (4)
This utility $u_{t}$ is the combination of all factors that the individual considers when making a decision. Here, we expressed the utility $u_{t}$ as a sum of the instantaneous stimulus $s_{t}$ and the internal memory $m_{t}$ to keep the utility calculation simple. However, other combinations of these two factors can also be considered. Likewise, the model can be extended to account for additional factors that are relevant to an individual’s decision-making process.
The individual chooses its next action $a_{t+1}$ at the next time step $t+1$ based on the utility $u_{t}$. The utility $u_{t}$ is passed to a logistic decision function:
$p_{a_{t+1}}=\frac{1}{1+e^{-k(u_{t}-b)}}$ (5)
@@
where the steepness $k$ indicates the individual’s certainty on the informational value of any given utility and $b$ is the midpoint of the function which we set to 0. The logistic decision function returns a probability $p_{a_{t+1}}$ indicating the individual’s likelihood of staying with the current action $a_{t}$ in the next time step $t+1$. Hence, the individual’s next action $a_{t+1}$ will be the same as the current action $a_{t}$ with probability $p_{a_{t+1}}$ or will be the other action with probability $1-p_{a_{t+1}}$.
The individual’s behavioural tendencies are characterised by two parameters. The first parameter is the memory decay $\tau$ which specifies the timescale of the individual’s memory and hence determines how much information the individual can recall from its past experiences. A higher $\tau$ indicates longer memory, while a smaller $\tau$ means shorter memory. The second parameter that characterises the individual’s behavioural tendencies is the logistic decision function’s steepness $k$ which indicates the individual’s certainty about the value of the information it has accumulated. Given a piece of information, an individual with higher $k$ is more certain about its value, while an individual with lower $k$ is less certain.
II.2 Relationship to COM-B model of behaviour change
In designing the components of individual behaviour in our microsimulation model, we used as a basis the components of the COM-B model, a behavioural change theory that has been used to design and implement real-world policies to drive positive changes in people’s health behaviour Michie et al., (2011). The COM-B model proposes that capability (C), opportunity (O), and motivation (M) are the three components that influence an individual’s behaviour (B). Capability refers to an individual’s psychological and physical abilities to enact a behaviour. Opportunity refers to the external factors that make it possible to enact the behaviour. Motivation refers to the cognitive processes involved in enacting this behaviour. Each component of our microsimulation model was motivated to provide a minimalistic representation of these three COM-B components.
In any computational model, there has to be an explicit mathematical formulation of how an individual makes decisions for behaviour to be specified. Hence, the individual must have a decision function. Without this decision function, then there is no behaviour. Here, we used a logistic function as the decision function, which by design, naturally represents the individual’s capability (C) to choose between two actions. While other formulations of the decision function can be used, the logistic decision function provides a simple mechanism for incorporating the capability of an individual to choose an action based on the available information (in the form of the utility). It also allows an infinite span of behaviours to be specified based on how the logistic decision function’s steepness $k$ is selected, where a higher $k$ indicates that an individual is more capable of discerning the value of information and a lower $k$ means that the individual is less capable. Hence, the logistic decision function not only conceptualises the capability of an individual in principle but also characterises the degree of this capability. The utility $u_{t}$ passed to this logistic decision function represents a form of motivation (M). This utility captures the combined effect of instantaneous reward signals (via the stimulus $s_{t}$) and reflective evaluations of past actions (via the memory $m_{t}$). With higher utility values, the individual becomes more motivated to choose one action over the other. Opportunity (O) is represented in the experimental setup where the reversal of reward signals provides an opportunity for the individual to change its behaviour, as mediated by the environment. Should barriers to learning this desired behaviour be included in the experiment, the way the environment provides the stimulus $s_{t}$ can be changed or the logistic decision function’s midpoint $b$ can be biased to make it easier or harder for an individual to change its behaviour.
II.3 Reversal learning experiments